OpenStudy (anonymous):

The function p(d)=1+ d/33 gives the pressure, in atmospheres (atm), at a depth d in the sea (d is in feet). Note that p(0)=1 atm,p(33)=2, and so on. Find the pressure at 20 feet. This one is really hard and I would be totally geeked if someone could answer this one.

OpenStudy (saifoo.khan):

u call this "REALLY HARD"?? :/

OpenStudy (anonymous):

Nobody can figure this out?

OpenStudy (anonymous):

\[p(20) = 1 + \frac{20}{33} = \frac{53}{33} \approx 1.60606.\]

OpenStudy (saifoo.khan):

lol, this is wasy!

OpenStudy (saifoo.khan):

easy*

OpenStudy (anonymous):

p(20) = 1 + (20/33) p(20) = 53/33

OpenStudy (saifoo.khan):

both are correct. ^^

OpenStudy (gw2011):

p(20)=1+20/33 p(20)=53/33=1.606